# Pushing the Accuracy Limit of Foundation Neural Network Models with Quantum Monte Carlo Forces and Path Integrals

### v2

@@ Abstract
We propose an end-to-end integrated strategy to produce highly accurate quantum chemistry (QC) synthetic datasets (energies and forces) aimed at deriving Foundation Machine Learning models for molecular simulation. Starting from Density Functional Theory (DFT), a ”Jacob’s Ladder” approach leverages computationally-optimized layers of massively parallel GPU-accelerated software with increasing accuracy. Thanks to Exascale, this is the first time that the computationally intensive calculation of Diffusion Quantum Monte Carlo (QMC) forces, and the combination of multi-determinant QMC energies and forces with selected-CI wavefunctions, are computed at such scale at the complete basis-set-limit. To bridge the gap between accurate QC and condensed-phase molecular dynamics, we leverage transfer learning to improve the FeNNix-Bio1 DFT-based foundation model. The resulting approach is coupled to path integrals adaptive sampling quantum dynamics to perform nanosecond reactive simulations at unprecedented accuracy. These results demonstrate the promise of Exascale to deepen our understanding of the inner machinery of complex biosystems.
@@ I Justification for Prize
  • Record Quantum Monte Carlo (Diffusion Monte Carlo - DMC) and Multideterminant DMC (from selected CI nodes (sCI)) computations were performed, including 20,338 QMC (energies-only), 20,251 QMC (energies + forces), and 2,000 multideterminant Diffusion Monte Carlo (energies + forces) calculations using determinants from sCI. These achieved double-precision performance of 124, 62, and 52 PFLOPS, respectively, on 2,048 nodes of Aurora (12,288 GPUs), corresponding to approximately 60%, 30%, and 25% of the system’s theoretical peak performance for energies-only, forces, and multideterminant forces evaluations. All QMC simulations were performed in real space and are thus free from basis-set errors.
  • For a 32-million-particule (1-million atom X 32 beads) Satellite Tobacco Virus system (STMV) Ring Polymer quantum molecular dynamics simulation with a foundation machine learning (ML) model (using the present project high accuracy quantum chemistry data) while using multi-replica via an unsupervised adaptive sampling strategy, time-to-solution of 0.44ns/day (4 replica, 128 nodes, 1024 GPUs, mixed precision on Jean Zay or 15 replica, 480 nodes, 5760 GPUs, mixed precision on Aurora). Establish new state-of-the-art for reactive foundation model atomistic quantum dynamics simulation of biological systems.
@@ II Performance Attributes
Type of method used (1) Quantum Chemistry
Type of method used (2) ML potential molecular dynamics
@@ III Overview of the problem
Simulating the structural dynamics of biological systems at full quantum accuracy is a daunting task requiring precise modeling of the Born-Oppenheimer electronic potential energy surface while including nuclear quantum effects in the dynamics to accurately reproduce condensed phase properties. Up to now, achieving such high-fidelity simulations has been prevented by the computational cost of first-principles quantum approaches that can only reach the required accuracy on very small systems and remains therefore a Graal for the molecular simulation community. In this context, neural networks potentials have started to gain speed and precision offering a clear alternative to ab initio molecular dynamics (AIMD) and recently even to force fields simulations thanks to high performance computing. Thus, the field evolves very fast, and larger foundation models dedicated to atomistic Molecular Simulation have been recently introduced. These single general-purpose ML potentials have various capabilities and can be applied to various types of applications, efficiently replacing quantum mechanical approaches. However, in practice, such ML approaches cannot perform computations at an accuracy better than the models’ reference data. Therefore, foundation models rely heavily on the quality and the quantity of their reference synthetic data. The choice of the quantum chemistry approaches and their capability to approximate the Schrodinger equation therefore have a considerable impact on the performance of the model. However, in principle, with enough extremely accurate small molecules quantum chemistry data, one could transfer such accuracy to very large systems. Additionally, on the molecular dynamics side, since foundation models can be seen as Born-Oppenheimer potential surfaces generators, it is critical to go beyond classical dynamics. Indeed, in this precise context where no ”condensed-phase parametrization” occurs like in usual force fields, it becomes paramount to include the effects of nuclei via quantum dynamics in order to predict accurate condensed-phase properties. Overall, the rise of exascale computing offers the possibility of designing a new, multi-level, pipeline for the generation of accurate quantum chemistry data that would strongly impact future foundation machine learning model developments. Indeed, ultimately, these latter could become capable of performing large scale, chemically accurate, molecular simulations.
@@ IV-1 Quantum Chemistry Datasets
In that context, the present consensus for the computation of reference quantum chemistry datasets is to use Density Functional Theory (DFT) potential energy surfaces which provide a reasonable accuracy at an affordable cost. However DFT’s precision strongly depends on the choice of the functional, which is by essence non-exact and therefore not fully transferable across the chemical space, especially if one wants to simultaneously model structures, weak interactions and reactivity of complex biological systems. Therefore, there is a need for more exact ab initio methods. for several decades CCSD(T)[74, 82] – often termed the ”gold standard” – has been widely employed. Nonetheless, CCSD(T) exhibits steep $O(N^{7})$ scaling, and its local approximations (e.g., DLPNO-CCSD(T)) rely on domain truncations that are typically benchmarked against known data but can fail if the chosen domains omit critical electron correlation[82]. As a result, such local-correlation variants, while computationally attractive for large systems, can yield unpredictable errors outside the familiar chemical space. Similarly, Density Matrix Renormalization Group (DMRG) offers high accuracy for strongly correlated systems but remains challenging for broad chemical data generation due to the need for carefully chosen active spaces.[110]
@@
Unlike traditional quantum chemistry methods, which rely on finite basis sets and must extrapolate to the complete basis set (CBS) limit or depend on error cancellation and ad hoc corrections, QMC—particularly DMC—operates directly in real space. As a result, it is inherently free from basis set incompleteness error and provides results effectively in the infinite-basis limit without the need for basis-set tricks or extrapolation schemes.
Generating multideterminant expansions to reduce nodal-surface errors has proven challenging. Methods such as CASSCF have been recognized for decades as effective but require extensive prior knowledge of the target system’s chemical properties, which limits their suitability for broad machine-learning applications. More recently, Caffarel and co-workers[32] revived the concept of selected Configuration Interaction (sCI)[38], automating the construction of multideterminant expansions in a largely black-box manner and enabling various sCI flavors[96, 37]. As a multideterminant trial wavefunction, sCI systematically improves nodal accuracy. Coupled with an optimized Jastrow factor, it resolves both static and dynamic correlations by addressing strong-correlation effects through the sCI expansion while capturing dynamical correlations in the Jastrow[12, 77, 53, 52, 55, 78, 54, 65, 67, 29]. In turn, QMC forces can validate or directly compute sCI forces, establishing a synergistic approach with high reliability[89].
Nevertheless, the computational cost of these methods remains substantial, restricting routine usage for many years. Over the past two decades, the community has extensively developed QMCPACK[45, 43, 61, 60] to exploit leadership-class supercomputers—from IBM Blue Gene architectures to exascale systems such as Aurora and Frontier at Argonne and Oak Ridge National Laboratories. Early efforts prioritized a GPU-based energy estimator, particularly via efficient B-spline orbital evaluations on massively parallel platforms, targeting material sciences application. Building on this foundation, the present work introduces new optimizations and data layouts that optimize the LCAO branch of QMCPACK for GPU acceleration for molecular applications.
@@ IV-A Foundation Machine Learning Models for atomistic molecular dynamics Simulations
The application of Neural Networks to molecular dynamics is recent and started with the availability of suitable neural architecture  [9, 6]. Since then, numerous neural network potentials (NNPs) have been proposed [20, 14, 9, 83, 90, 80, 79, 31, 56, 108, 104, 73, 97, 30, 8, 64, 34, 35, 21, 109, 68, 46, 103] while their high performance implementations reached a mature stage [88, 41, 39] and as specialized libraries are now available to NNPs developers  [81, 22, 105, 69]. Very recently, in link with the breakthroughs of the AlphaFold [42] and RoseTTAFold [5] protein structure foundation models, a major shift in the developments started to reshape molecular simulation with the introduction of the MACE-MP-0 foundation model for molecular dynamics. [7]. As explained by its authors, MACE is a single general-purpose ML potential, designed for material science atomistic simulations. Its capabilities range from single molecule to full materials simulations, and enables modeling of chemical reactivity. Beyond solids, its extended transferability allowed it to be tested even on small proteins where stable molecular dynamics could be obtained. However, since MACE and its associated neural architecture are primarily designed for materials, its applicability to condensed phase simulations of large complex biological molecules remains limited both in term of accuracy and computational performances. For these reasons, some of us recently introduced the FeNNix-Bio1 Foundation Model[72] for atomistic molecular dynamics simulations. Note that its associated paper can be seen as the machine learning counterpart of the present Exascale computing report as both projects were launched together and the FeNNix-Bio1 development used some preliminary data extracted from the present quantum chemistry pipeline. Thus, it is able to perform atomistic condensed phase molecular dynamics simulations of biological systems while allowing for reactivity modeling. FeNNix-Bio1 is fast and accurate, being able to address simultaneously the issues of the computational requirements generated by the biological timescales and the difficult transferability of charged-species, that are numerous in Chemistry and Biology. Finally, FeNNix-Bio1 is also able to compute meaningful condensed-phase properties since it explicitly takes into account nuclear quantum effects in quantum dynamics. [18, 62, 70]. This allows the foundation model to predict accurate condensed phase properties such as free energies, solvent properties, etc. .
@@ V-A Dataset Generation: High Performance Quantum Chemistry Pipeline
We first introduce our GPU-accelerated quantum chemistry database generation Exascale protocol which has been optimized to able to produce the required energies and forces. Organized as a “Jacob’s Ladder” approach (see Figure  1) with methods of increasing complexity, it leverages 3 different layers of computationally-optimized massively parallel high performance softwares dealing with increasing levels of accuracy. Note that exact, i.e. Full-Configuration interaction (or Full-CI) computations in the complete basis set limit, numerically reaching the quality of the electronic time-independent, non-relativistic, Schrödinger’s equation quality computations are mostly out of reach of classical computers due to the existence of a complexity exponential wall (see[98] and references therein). Such a quality may be reached in a near future using quantum computers and fault tolerant Quantum Processing Units (QPU)[106]. For now, and since the goal of our work is to apply the quantum data to chemistry and biology simulations, the multi-level GPU-accelerated computational strategy proposed here aims at reaching the chemical accuracy which is the required level of precision needed to perform realistic chemical, i.e. comparable to experiment, predictions. Chemical accuracy implies errors below 1 kcal/mol (1.59 mH/particle) with respect to the exact Schrödinger total energy. Since forces are the cornerstone of the design of machine learning potentials, each proposed level includes the efficient evaluation of the energy gradients.
@@ V-A1 The SPICE2(+)-ccECP dataset
In order to test the computational pipeline, we computed a first training set denoted SPICE2(+)-ccECP. It used the molecular configurations of the SPICE2 database[24, 91], which we extended with a carefully selected subset of 100k configurations from the ANI-2X dataset[21]. These latter were selected via a round of active learning using the uncertainty given by a preliminary version of the FeNNix-Bio1 model and recomputed using our selected level of theory.This combined dataset provided comprehensive coverage of the conformational space relevant to biomolecular systems, including diverse functional groups, a wide range of environmental conditions, and various non-equilibrium configurations critical for modeling reactive processes. Diverging from the original SPICE2, we combined in the present work DFT with the correlation-consistent effective core potentials (ccECP)[100] and their companion aug-cc-pVTZ basis sets[100], as implemented in the PySCF package[94]. The choice of the present effective core potentials conserves accuracy while increasing the computational efficiency compared to the explicit inclusion of core electrons. Moreover, this enables us to extend the SPICE2 dataset to post Hartree-Fock methods such as QMC where the use of valence-only calculations is crucial. Overall, the dataset encompasses 2,108,628 conformations that were all treated (energies and forces) at the DFT level. Among these latter, 20,251 conformations were computed using QMC (energies and forces), 20,338 QMC (energies-only) and 2000 (energies and forces) with multideterminant QMC with selected CI wavefunctions.
@@ V-A2 Density Functional Theory
For the first layer of our quantum chemistry database generation protocol, we performed Density Functional Theory (DFT) calculations using the PySCF code[93, 51, 102], specifically leveraging the GPU-accelerated implementation. We strategically allocated computational resources based on system size. For molecules up to 175 electrons (approximately 1.3 million configurations), calculations were executed on the Leonardo supercomputer at Cineca (Italy), larger molecular systems with up to 275 electrons (approximately 0.8 million configurations) were computed on the Polaris supercomputer at Argonne National Laboratory. Both supercomputers feature NVIDIA A100 GPUs-40Gb. The largest molecules with more than 275 electrons (65,521 conformations) were computed on Aurora Supercomputer which features Intel Xe GPUs. A significant challenge for large-scale DFT calculations is the computation and storage of electron repulsion integrals (ERIs), which cannot be generated on-the-fly on GPUs for larger molecules due to memory limitations of the A100-40Gb cards. The conventional CPU implementation would require over 6 hours per molecule on AMD Epyc 7303 processors and up to 300GB of temporary storage per calculation, making the complete dataset generation computationally infeasible. To overcome these limitations, we employed our gpu4mrh implementation, which optimizes memory usage and enables efficient offloading of integral calculations to GPUs even for large molecular systems.
@@
The DFT calculations were distributed across the Nvidia A100 nodes systems (Polaris and Leonardo) with one molecule assigned per GPU, using 4 GPUs per node. This distribution strategy optimized the trade-off between parallel efficiency and memory requirements. Specifically, the DFT calculations achieved approximately 16 TFLOPS per node (40% of theoretical peak) for smaller molecules (0–100 electrons), 20 TFLOPS per node (50% of theoretical peak) for intermediate-sized molecules (100–200 electrons), and 28 TFLOPS per node (70% of theoretical peak) for larger molecules (200–250+ electrons). Each DFT simulation was distributed across 128 nodes, with 4 Nvidia A100 GPUs per node (512 GPUs per simulation in total). This large-scale deployment was essential, as it was the only feasible way to generate the extensive dataset of 2.1 million molecular configurations within the allocated time frame. Specifically, the simulations achieved peak performances of approximately:
@@
On Aurora, we used PySCF coupled with the gpu4mrh code that leverages the Intel Xe architecture through efficient data layout patterns and kernel optimizations to maximize use of matrix-multiply operations central to quantum chemistry calculations. The gpu4mrh code, initially developed to accelerate multi-configurational electronic structure methods, uses a lightweight Python/C++ abstraction layer to seamlessly offload fundamental operations from PySCF (namely, formation of the Coulomb and exchange matrices) to GPUs via CUDA, HIP, or SYCL and respective vendor-optimized math libraries (e.g. MKL on Aurora). With an efficient hashing strategy to minimize data-transfer GPU-host data transfers, the majority of the GPU runtime is spent computing single and batched matrix-matrix products concurrently across multiple GPUs in a compute node. The gpu4mrh code was used as a drop-in replacement enabling SCF energy calculations to leverage the full 768 GB of HBM from the 6 GPUs on each Aurora node. The cost of this implementation is dominated by several dense gemm operations, so it is able to make use of vendor-optimized BLAS routines and achieves high GPU utilization for the systems considered here. In total, we performed calculations on 2,1 million molecular configurations, requiring approximately 20k nodes-hours on Leonardo, 35k node-hours on Polaris and 25k nodes hours on Aurora. The resulting database of energies and forces serves as the foundation layer for our multi-level quantum chemistry approach.
In addition to providing the first level of our Jacob’s Ladder approach, these DFT calculations served as essential inputs for our higher-level quantum chemistry methods. The DFT wavefunctions were used as trial wavefunctions for the single-determinant Quantum Monte Carlo calculations. Additionally, the one- and two-electron integrals generated during the DFT calculations were utilized as inputs for the selected Configuration Interaction (sCI) calculations, ensuring consistency across all three levels of theory.

### v3

@@ Abstract
We propose an end-to-end integrated strategy to produce highly accurate quantum chemistry (QC) synthetic datasets (energies and forces) aimed at deriving Foundation Machine Learning models for molecular simulation. Starting from Density Functional Theory (DFT), a ”Jacob’s Ladder” approach leverages computationally-optimized layers of massively GPU-accelerated software with increasing accuracy. Thanks to Exascale, this is the first time that the computationally intensive calculation of Quantum Monte Carlo forces (QMC), and the combination of multi-determinant QMC energies and forces with selected-Configuration Interaction wavefunctions, are computed at such scale at the complete basis-set limit. To bridge the gap between accurate QC and condensed-phase Molecular Dynamics, we leverage transfer learning to improve the DFT-based FeNNix-Bio1 foundation model. The resulting approach is coupled to path integrals adaptive sampling quantum dynamics to perform nanosecond reactive simulations at unprecedented accuracy. These results demonstrate the promise of Exascale to deepen our understanding of the inner machinery of complex biosystems.
@@ I Justification for Prize
  • Record Diffusion Quantum Monte Carlo (DMC): energies-only, energies/forces and Multideterminant DMC: energies/forces; double-precision performance: 124, 62, 52 PFLOPS; 60%, 30%, 25% Aurora’s theoretical peak; 12,288 GPUs.
  • Record time-to-solution for full quantum dynamics foundation machine learning model simulation. Time-to-solution: 0.44ns/day mixed-precision (1024 GPUs, Jean Zay; 5760 GPUs Aurora).
@@ II Performance Attributes
Type of methods used Quantum Chemistry, ML potentials MD
@@ III Overview of the problem
Simulating the structural dynamics of biological systems at full quantum accuracy is a daunting task requiring the precise modeling of the Born-Oppenheimer (BO) electronic potential energy surface[1] while including nuclear quantum effects (NQEs) in the dynamics to accurately reproduce condensed phase properties[2]. Up to now, achieving such high-fidelity simulations has been prevented by the computational cost of first-principles quantum approaches that can only reach maximum accuracy on the BO surface of small few-atom systems[1]. Full quantum accuracy MD remains therefore a Graal for the molecular simulation community. In this context, neural networks potentials[3] have started to gain speed and precision offering a clear alternative to ab initio molecular dynamics (AIMD)[4] and recently even to force fields simulations[5] thanks to GPU-accelerated high performance supercomputing[6]. Thus, the field evolves very fast, and larger general-purpose foundation neural network models dedicated to atomistic Molecular Simulation have been recently introduced[7]. However, these ML approaches cannot perform computations at an accuracy better than their reference data. Consequently, the choice of the synthetic quantum chemistry approaches used for the reference datasets and their capability to approximate the Schrödinger equation have a considerable impact. Additionally, once reaching high precision on the BO energies, it becomes paramount to accurately reproduce experimental observations. To take full advantage of the high accuracy achieved by neural networks, one must take into account for NQEs explicitly[2, 8] via some form of quantum dynamics. This is mandatory for the accurate prediction of condensed-phase properties. Overall, the rise of exascale computing offers the possibility of designing a new, multi-level pipeline for the generation of accurate quantum chemistry datasets. This will enable to derive ML models capable of performing large scale quantum dynamics molecular simulations towards accurately predicting the structural dynamics and chemical reactivity of highly complex biosystems.
@@ IV-1 Quantum Chemistry Datasets
The present consensus for the computation of reference quantum chemistry datasets is to use DFT potential energy surfaces which provide a reasonable accuracy at an affordable cost. However DFT’s precision strongly depends on the choice of the functional, which is by essence non-exact and therefore not fully transferable across the chemical space. Therefore, there is a need for more exact ab initio methods. For several decades CCSD(T)[9, 10] often termed the ”gold standard”, has been widely employed. Nonetheless, CCSD(T) exhibits steep $O(N^{7})$ scaling, and its local approximations (e.g., DLPNO-CCSD(T)) rely on domain truncations that are typically benchmarked against known data but can fail if the chosen domains omit critical electron correlation[10]. As a result, such local-correlation variants, while computationally attractive for large systems, can yield unpredictable errors outside the familiar chemical space. Similarly, Density Matrix Renormalization Group (DMRG) offers high accuracy for strongly correlated systems but remains challenging for broad chemical data generation due to the need for carefully chosen active spaces.[11]
@@
Unlike traditional quantum chemistry methods, which rely on finite basis sets and must extrapolate to the complete basis set (CBS) limit or depend on error cancellation and ad hoc corrections, QMC, and particularly DMC, operates directly in real space. As a result, it is inherently free from basis set incompleteness error and provides results effectively in the CBS limit without the need for basis-set tricks or extrapolation schemes.
Generating multideterminant expansions to reduce nodal-surface errors has proven to be challenging. Methods such as CASSCF have been recognized for decades as effective but require extensive prior knowledge of the target system’s chemical properties, which limits their suitability for broad machine-learning applications. More recently, Caffarel and co-workers[31] revived the concept of selected Configuration Interaction (sCI)[32], automating the construction of multideterminant expansions in a largely black-box manner and enabling various sCI flavors[33, 34]. As a multideterminant trial wavefunction, sCI systematically improves nodal accuracy. Coupled with an optimized Jastrow factor, it resolves both static and dynamic correlations by addressing strong-correlation effects through the sCI expansion while capturing dynamical correlations in the Jastrow[25, 35, 36, 37, 38, 39, 40, 41, 42, 43]. In turn, QMC forces can validate or directly compute sCI forces, establishing a synergistic approach with high reliability[44].
Nevertheless, the computational cost of these methods remains substantial, restricting routine usage for many years. Over the past two decades, the community has extensively developed QMCPACK[45, 46, 47, 48] to exploit leadership-class supercomputers—from IBM Blue Gene architectures to exascale systems such as Aurora and Frontier at Argonne and Oak Ridge National Laboratories respectively. Early efforts prioritized a GPU-based energy estimator, particularly via efficient B-spline orbital evaluations on massively parallel platforms, targeting material sciences application. Building on this foundation, the present work introduces new optimizations and data layouts that optimize the LCAO branch of QMCPACK for GPU acceleration for molecular applications.
@@ IV-A Foundation Machine Learning Models for atomistic molecular dynamics Simulations
The application of Neural Networks to MD is recent and started with the availability of suitable neural architectures  [49, 50]. Since then, numerous neural network potentials (NNPs) have been proposed [51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 3]. As NNPs’ high performance implementations reached a mature stage [73, 6, 74], specialized libraries  [75, 76, 77, 78] are now available to developers facilitating the design of new NNPs. Very recently, following the advances of the AlphaFold [79] and RoseTTAFold [80] protein structure foundation models, a shift in the developments started with the introduction of the MACE-MP-0 foundation model for MD. [7]. MACE is a single general-purpose ML potential, designed for material science atomistic simulations. Its capabilities range from single molecule to full materials simulations, and enables modeling of chemical reactivity. However, since MACE and its associated neural architecture are primarily designed for materials, its applicability to condensed phase simulations of large complex biological molecules remains limited both in term of accuracy and computational performances. For these reasons, some of us recently introduced the FeNNix-Bio1 Foundation Model[81] for atomistic MD simulations going beyond the initial FeNNix equivariant NNP[82]. Thus, FeNNIx-Bio1 is able to perform reactive atomistic condensed phase MD simulations of biological systems addressing simultaneously the issues of the computational requirements generated by the biological timescales and the difficult transferability of charged-species. Finally, it is also able to compute meaningful condensed-phase properties (free energies, solvent properties,…) as it takes explicitly into account NQEs [83, 8, 84].
@@ V-A Dataset Generation: High Performance Quantum Chemistry Pipeline
We first introduce our GPU-accelerated quantum chemistry database generation Exascale protocol which has been optimized to able to produce the required energies and forces. Organized as a “Jacob’s Ladder” approach (see Figure  1) with methods of increasing complexity, it leverages 3 different layers of computationally-optimized massively parallel high performance softwares dealing with increasing levels of accuracy. Note that exact, i.e. Full-Configuration interaction (or Full-CI) computations in the CBS limit, numerically reaching the quality of the electronic time-independent, non-relativistic, Schrödinger’s equation quality computations are mostly out of reach of classical computers due to the existence of a complexity Exponential Wall problem[85, 86]. The multi-level GPU-accelerated computational strategy proposed here aims at reaching the chemical accuracy which is the required level of precision needed to perform realistic chemical, i.e. comparable to experiment, predictions. It implies errors below 1 kcal/mol (1.59 mH/particle) with respect to the exact Schrödinger total energy. Since forces are the cornerstone of the design of ML potentials, each proposed level includes the efficient evaluation of the energy gradients.
@@ V-A1 The SPICE2(+)-ccECP dataset
Testing the computational pipeline, we computed a first training set denoted SPICE2(+)-ccECP. It used the molecular configurations of the SPICE2 database[87, 88], which we extended with a carefully selected subset of 100k configurations from the ANI-2X dataset[68]. These latter were selected via a round of active learning using the uncertainty given by a preliminary version of the FeNNix-Bio1 model and recomputed using our selected level of theory. This combined dataset provided comprehensive coverage of the conformational space relevant to biomolecular systems, including diverse functional groups, a wide range of environmental conditions, and various non-equilibrium configurations critical for modeling reactive processes. Diverging from the original SPICE2, we combined in the present DFT work with correlation-consistent effective core potentials (ccECP)[89] and their companion aug-cc-pVTZ basis sets[89], as implemented in the PySCF package[90]. This choice conserves accuracy while increasing the computational efficiency compared to the explicit inclusion of core electrons while enableing us to extend the SPICE2 dataset to post Hartree-Fock methods such as QMC where the use of valence-only calculations is crucial. The dataset encompasses 2,108,628 conformations that were all treated (energies and forces) at the DFT level. Among these latter, 20,251 conformations were computed using QMC (energies and forces), 20,338 QMC (energies-only) and 2000 (energies and forces) with multideterminant QMC with selected CI wavefunctions.
@@ V-A2 Density Functional Theory
For the first layer of our quantum chemistry database generation protocol, we performed DFT calculations using the PySCF code[91, 92, 93], specifically leveraging the GPU-accelerated implementation. We strategically allocated computational resources based on system size. For molecules up to 175 electrons (approximately 1.3 million configurations), calculations were executed on the Leonardo supercomputer at Cineca (Italy), larger molecular systems with up to 275 electrons (approximately 0.8 million configurations) were computed on the Polaris supercomputer at Argonne National Laboratory. Both supercomputers feature NVIDIA A100 GPUs-40Gb. The largest molecules with more than 275 electrons (65,521 conformations) were computed on Aurora Supercomputer which features Intel Xe GPUs. A significant challenge for large-scale DFT calculations is the computation and storage of electron repulsion integrals (ERIs), which cannot be generated on-the-fly on GPUs for larger molecules due to memory limitations of the A100-40Gb cards. The conventional CPU implementation would require over 6 hours per molecule on AMD Epyc 7303 processors and up to 300GB of temporary storage per calculation, making the complete dataset generation computationally infeasible. To overcome these limitations, we employed our gpu4mrh implementation, which optimizes memory usage and enables efficient offloading of integral calculations to GPUs even for large molecular systems.
@@
The DFT calculations were distributed across the Nvidia A100 nodes systems (Polaris/Leonardo) with one molecule assigned per GPU, using 4 GPUs per node. This distribution strategy optimized the trade-off between parallel efficiency and memory requirements. Specifically, the DFT calculations achieved approximately 16 TFLOPS per node (40% of theoretical peak) for smaller molecules (0–100 electrons), 20 TFLOPS per node (50% of theoretical peak) for intermediate-sized molecules (100–200 electrons), and 28 TFLOPS per node (70% of theoretical peak) for larger molecules (200–250+ electrons). Each DFT simulation was distributed across 128 nodes, with 4 Nvidia A100 GPUs per node (512 GPUs per simulation in total). This large-scale deployment was essential, as it was the only feasible strategy to generate the extensive dataset of 2.1 million molecular configurations within the allocated time frame. Specifically, the simulations achieved peak performances of approximately:
@@
On Aurora, we used PySCF coupled with the gpu4mrh code that leverages the Intel Xe architecture through efficient data layout patterns and kernel optimizations to maximize use of matrix-multiply operations central to quantum chemistry calculations. The gpu4mrh code, initially developed to accelerate multi-configurational electronic structure methods, uses a lightweight Python/C++ abstraction layer to seamlessly offload fundamental operations from PySCF (namely, formation of the Coulomb and exchange matrices) to GPUs via CUDA, HIP, or SYCL and respective vendor-optimized math libraries (e.g. MKL on Aurora). With an efficient hashing strategy to minimize data-transfer GPU-host data transfers, the majority of the GPU runtime is spent computing single and batched matrix-matrix products concurrently across multiple GPUs in a compute node. The gpu4mrh code was used as a drop-in replacement enabling SCF energy calculations to leverage the full 768 GB of HBM from the 6 GPUs on each Aurora node. The implementation cost is dominated by several dense gemm operations, so it is able to make use of vendor-optimized BLAS routines and achieves high GPU utilization for the systems considered here.
In total, we performed computed 2,1 million molecular configurations, requiring approximately 20k nodes-hours on Leonardo, 35k node-hours on Polaris and 25k nodes hours on Aurora. The resulting database of energies and forces serves is the foundation layer for our multi-level quantum chemistry approach.
In addition to providing the first level of our Jacob’s Ladder approach, these DFT calculations served as essential inputs for our higher-level quantum chemistry methods. The DFT wavefunctions were used as trial wavefunctions for the single-determinant QMC calculations. Additionally, the one- and two-electron integrals generated during the DFT calculations were utilized as inputs for the selected Configuration Interaction (sCI) calculations, ensuring consistency across all three levels of theory.